In a large population of randomly reproducing rabbits, a recessive allele r comprises 80% of the alleles for a gene. What percentage of the rabbits would you expect to have the dominant phenotype?

  • 36%

  • 32%

  • 40%

  • 50%

  • 64%
Answer #1

Option a is correct. 36% of the rabbits would be expected to have the dominant phenotype.

This is because recessive allele comprises 80% of the population so the frequency of recessive allele is 0.8

The frequency of dominant allele will be 1 - 0.8 = 0.2

According to Hardy weinberg equilibrium

p^2 + 2pq + q^2 = 1

The percentage of rabbits to dominant phenotype is p^2 + 2pq

That is 0.2^2 + 2(0.2)(0.8)

= 0.4 + 0.32

= 0.36 or 36 %